Solve the differential equation y"=1+(y')^2
1
Expert's answer
2020-04-15T17:54:23-0400

Given "y''=1+y'^2."

Put "z = y'." Then the given equation becomes,

"z' = 1+ z^{2}" .

"\dfrac{dz}{dx}=1+z^2\\
\dfrac{dz}{1+z^2}=dx"

Integrating both sides, we get

"\arctan z = x + c_{1}\\
z = \tan (x+c_{1})\\
y'=\tan (x+c_{1})~~\text{(Using the substitution we have taken)}\\
\dfrac{dy}{dx}=\tan (x+c_{1})\\
\text{Integrating both sides with respect to $x$,}\\
y = \ln\sec (x+c_{1})+c_{2}"